what is 2x-3y=12
x=4y+1
substituion form

Found 2 solutions by checkley71, funmath:
Answer by checkley71(8403)

2X-3Y=12
X=4Y+1 MULTIPLY THIS EQUATION BY -2 AND ADD
-2X+8Y=-2
2X-3Y=12
--------------------------
5Y=10
Y=10/5
Y=2
X=4Y+1
X=4*2+1
X=8+1
X=9
PROOF
2*9-3*2=12
18-6=12
12=12
65/5=12

Answer by funmath(2933)

what is
L1) 2x-3y=12
L2) x=4y+1
substituion form
:
Substitute the right side of L2 into L1 for x and solve for y
2(4y+1)-3y=12
8y+2-3y=12
5y+2=12
5y+2-2=12-2
5y=10
5y/5=10/5
y=2
:
Substitute 2 in for y in L2 and solve for x
x=4(2)+1
x=8+1
x=9
The solution (x,y)=(9,2)
This is the point that both equations share. Garphically, its' the point of intersection of the two lines.
Happy Calculating!!!